OpenStudy (anonymous):

reduce to lowest terms- 738/882 srry i am horrible at this

OpenStudy (whynot):

-738/882=369/441=123/147=41/49

OpenStudy (anonymous):

is that negative? 369/441= 123/147= 41/49

OpenStudy (anonymous):

no